3) What is the molarity of a saturated solution of Nilo) (K.,-47 x 10)?(0 pts)
Answer #1

Answer:

Ni(IO3)2 (aq) \rightarrowNi2+ (aq) + 2 IO3- (aq).

Ksp = [Ni2+\[IO3-]2. .............. (1)

Let Solubility of Ni(IO3)2 (aq) = "S" M

So, [Ni2+] = "S" M & [IO3-] = "2S" M

Hence we write,

Ksp = (S)(2S)2.

Ksp = 4S3.

Ksp (Ni(IO3)2) = 4.7*10-5.

I.e. 4S3 = 4.7*10-5.

S3 = 4.7*10-5./4

S3 = 11.75*10-6.

S = 2.27*10-2 M

For saturated solution,(I.e. Salt ionizes completely)

[Ni(IO3)2] = S = 2.27*10-2 M.

Is the molarity of (Ni(IO3)2).

=========================XXXXXXXXXXXXXXXXXXXX====================